the marginal social cost of garbage disposed is: MSCd (gd) = 10, where gd is a 50 gallon drum of garbage disposed. The marginal cost of recycling is: MSCr (gr) = 2gr, where gr is a 50 gallon can of recyclables collected. The marginal benefit for the resident of garbage removed from the residence is: MB(gd, gr) = 800 - 2(gd + gr).

At an economically efficient level, how many [ ] 50-gallon cans of materials a resident shall recycle? (Assuming most of residential waste generated can be recycled, i.e., the resident can recycle as much as desired.)


1
Expert's answer
2021-10-20T16:41:52-0400

Given

Marginal social cost of disposable = 10

Marginal cost of recycling = 2gr

"Marginal \space Benefit = 800 - 2(gd+gr)"

Economic efficient level is achieved where Marginal benefit = Marginal cost

"800 - 2(gd+gr) =2gr\\

800 - 2(10 + gr) = 2gr\\

800 - 20 - 2gr = 2gr\\

780 = 4gr

gr = \frac{780}{4}\\

gr = 195"

therefore the optimum level recycling is 195